OpenStudy (anonymous):

What is the value of this expression if s = −2 and t = −5? 3(s − t) A. −21 B. −9 C. −1 D. 9

OpenStudy (anonymous):

just plug in the numbers: 3(s − t) =3(-2-(-5)) =3*3 =9

OpenStudy (anonymous):

Thank you.